OpenStudy (anonymous):

The adjacent sides of a parallelogram measure 14 centimeters and 20 centimeters, and one angle measure 57. What is the area of the parallelogram?

OpenStudy (anonymous):



OpenStudy (anonymous):

sin 57 = x/14 x = 14sin57

OpenStudy (anonymous):

x=11.7414 area = 11.7414 * 20 = 234.828